OpenStudy (anonymous):

Consider f(x) = x^3 - 2x^2 - 5x + k Find the value of k if (x+2) is a factor of f(x)

OpenStudy (anonymous):

k=6

OpenStudy (anonymous):

I figured that out but I don't know how to explain

OpenStudy (anonymous):

if x+2 is a factor then when u plug in -2 u should get 0 as a result so: 0=(-2)^3-2(-2)^2-5(-2)+k 0=-8-8+10+k 0=-16+10+k 0=-6+k k=6